OpenStudy (anonymous):

0.4x+7≤0.8x-2 solve. (simplify your answer; type an inequality symbol then type an integer or fraction)

OpenStudy (anonymous):

make your life easy and multiply by 10 to get \[4x+70\leq8x-20\]

OpenStudy (anonymous):

subtrac 4x from both sides, add 20 to both sides to get \[90\leq 4x\]

OpenStudy (anonymous):

then divide both sides by 4 and get \[\frac{90}{4}\leq x\]

OpenStudy (anonymous):

is that simplified?

OpenStudy (anonymous):

I'm sorry I simplified this and it still says that it is wrong.